Operations Report – 12 April 2018

SOL 4

Name of person filing report: Joe Dituri

Non-nominal systems: 1. The main hab door broke during the wind.
2. The circuit breaker tripped.

Notes on non-nominal systems: During severe winds again today, the door swung open HARD and almost came off its hinge. We’ve performed a sturdy and strong repair that is holding beautifully so far.

After it became very cold in the hab, we checked the circuit breaker and noticed it had tripped. It was reset.

Generator (hours run): 13 hours

Solar ‒ SOC 98%

Diesel – 75% fuel

Propane – 66% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) –

Water (trailer) – Empty, tank being cleaned in RAM.
Water (static) – 960 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133383.5

Toilet tank emptied: Yes. Last emptied April 11, 2018.

ATV’s Used: Blue ATVs #1, #2, and #3

Oil Added? No But checked and OK.

ATV Fuel Used: 10%

# Hours the ATVs were Used today: 2

Notes on ATVs: All working well and ready for use.

Deimos rover used: No
Hours: 119.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Spirit rover used: No
Hours: 31
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

Opportunity rover used: No
Hours: 29.8
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 25.2
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? No

General notes and comments: Crew is doing well, we had to stay on alert today for the strong winds. Will remain alert for continued winds tomorrow.

Summary of internet: 198MB at 1930.

Summary of suits and radios: Suits all charging, 4 one-piece suits and 1 two piece suit used today, radios working fine.

Summary of Hab operations: Repaired main hab door latch and wooden rail after strong winds blew the door all the way out, broke the latch, and knocked the rail off the stairs. Cleaned and decontaminated the trailer water tank in the RAM. The tank requires further cleaning.

Summary of GreenHab operations: See GreenHab report.

Summary of ScienceDome operations: None

Summary of RAM operations: The trailer water tank was cleaned in the RAM.